HUBBLE SPACE TELESCOPE DAILY REPORT #2269 PERIOD COVERED: 0000Z (UTC) 12/11/98 - 0000Z (UTC) 12/14/98 Daily Status Report as of 348/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Thirty-Seven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.2 Completed WF/PC-2 6844 (M87 Globular Clusters {#6746 Parallel Field Follow-up }) The WF/PC-2 was used to make observations of the globular clusters which are located on the "south" end of galaxy M-87. Sub-arcsecond dithering was employed to reduce the effect of undersampling of the point spread function. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Twenty-Two Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.4 Completed Nine Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.5 Completed Eleven Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.6 Completed Six Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Ten Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 6347 (A Search for Jets in Planetary Nebulae) The WF/PC-2 was used to make observations of the planetary nebulae (class PNe) PK190-17D1 in order to search for the presence of possible narrow jets. The current model for PNe formation indicates that jets are an unexpected phenomena, yet there is a growing body of circumstantial evidence that suggests narrow jets are common feature in PNe. The observations were completed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxies VCC899 and VCC888 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Two Sets of NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed NIC/2/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Cameras Two and Three) was used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Four Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.13 Completed Two Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.14 Completed Two Sets of FOC/96 6350 (FOC spectroscopy of BALQSOs) The Faint Object Camera (f/96) was used to make ultraviolet spectrographic images of BALQSO 2201-1834 and 2212-1759 which are known to show broad emission lines from abundant ions. This will help us improve our knowledge of these BAL class objects. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed STIS/CCD/MA1 7747 (The Relationship Between Lyman-alpha Clouds and Galaxies at Z<0.3) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of Lyman-Alpha clouds in galaxy/QSO PG0953+415. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed Three Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.17 Completed Six Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed Two Sets of WF/PC-2 6608 (Detailed Studies of Two Z > 3.5 Radio Galaxies) The WF/PC-2 was used to make moderate deep (i.e., six hours) observations of the radio galaxies Gal-0512+6 and Gal-0512+6 which are both located at a distance of Z > 3.5 . Both components appear to be embedded in a common halo of diffuse emission. The observations were completed as planned, and no problems were reported. 1.19 Completed STIS/CCD 7310 (The Interaction of Accretion Disks with Protostellar Binaries) The Space Telescope Imaging Spectrograph (CCD) was used to study the interaction of the accretion disk with the protostellar binary system LKHA332G2. This should determine how the accretion disk interacts with stars in young binary systems and will measure the rates that mass accumulates onto each component of the pair. The observations were executed as scheduled, and no anomalies were noted. 1.20 Completed STIS/CCD/MA2 8069 (Throughput of F25ND5 Filter) The Space Telescope Imaging Spectrograph (CCD/MA2) was used to make observations of star BD+75D325 in order to measure the throughput of the F25ND5 filter. Previous on orbit measurements of the throughput of this filter indicate that is about a factor of ~10 less efficient than the pre-launch estimates. The observations were completed as scheduled, and no anomalies were reported. 1.21 Completed WF/PC-2 6581 (Morphology and Photometry of Galaxies in Optically Selected High Redshift Clusters) The WF/PC-2 was used to make observations of the galactic cluster Gal-Clus-0317 in order to make photometric observations and to study its morphology. The high angular resolution imagery of HST, combined with our extensive ground based u, v, B, g, V, R, I and z photometry and spectroscopy, will allow the measurement of the morphological fractions in distant clusters as functions of redshift, color, spectral features, galaxy density, and global cluster properties such as core radius and ellipticity. The observations were executed as scheduled, and no anomalies were reported. 1.22 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxy 3C142.1. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.23 Completed STIS/CCD/MA1/MA2 7305 (The Physical Characteristics of Small- Scale ISM Structure) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the star HD6456 in order to measure the physical characteristics of small-scale ISM structure. A wavelength calibration observation was also taken. The observations were completed as planned, and no anomalies were reported. 1.24 Completed WF/PC-2 6363 (The Environments and Host Galaxies of BL-Lac Objects) The WF/PC-2 was used to make snapshot survey observations of the BL-Lac object PKS2201+04 in order to study its environment and host galaxy. The goal is to investigate systematic trends in host galaxy type, unusual morphological features, nearby companion galaxies, etc. The observations were executed as scheduled, and no anomalies were noted. 1.25 Completed WF/PC-2 6478 (The Evolution of Elliptical Galaxies in Distant Clusters) The WF/PC-2 was used to make observations in the V, R and I bands of the elliptical galaxies which are located in the distant galactic cluster Gal-Clus-0952+44. This will allow us to investigate the evolution of the cluster's ellipticals with redshift. The observations were completed as planned, and no problems were reported. 1.26 Completed WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 20 Successful : 20 Scheduled Reacquisitions : 26 Successful : 26 2.2 FHST Updates: Scheduled : 52 Successful : 52 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared fourteen times. During the weekend, several temperature parameters briefly flagged out of limits high: ICDZ04MN, EAXLK1T, ETANLL1T, and ESHLF21T. HSTARs were written for each occurrence. The NSSC-1 status buffer was dumped and reset at 346/1023Z per ROP NS-3. Forty-nine minutes of 32K engineering data was lost, beginning at 346/145930Z, due to an SSR misconfiguration -- a science playback was in progress when the engineering data was commanded to begin. HSTAR 6810 was written. As described in HSTAR 6812, the NICMOS dewar temperature (NDWTMP11) flagged out of limits high for 30 seconds beginning at 347/054859Z. As directed by ROP NS-5, an SI C&DH error was cleared at 347/1736Z. Per HSTAR 6814, a NICMOS status buffer message was received at 348/015604Z when an error bit was set while sending TPG command data. At 348/015628Z, another message was received stating that the timing pattern ID comparison had failed to match. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH